The pH of an acid solution is 6.00. Calculate the Ka for the monoprotic acid. The initial acid concentration is 0.010 M.

Ka =
Answer #1

HA <-----------> H+ + A-

pH = 6.00

[H+] = 10-pH = 10-6.00 = 1.0 x 10-6 M

at equilibrium [H+] = [A-] = 1.0 x 10-6 M

[HA] = 0.010 - 1.0 x 10-6 = 0.009999 M

Ka = [H+] [A-] / [HA]

Ka = [1.0 x 10-6] [1.0 x 10-6] / [0.009999]

Ka = 1.00 x 10-10